(a) The axes of two Polariods are placed at an angle so that the intensity of the incident unpolarized light is reduced to 1/4. What is that angle?
°

(b) If unpolarized light passes through two Polariods whose axes are placed at an angle of30 degree, how much would the intensity of the light be reduced to?
Answer #1

(a) from malus law I = (Io/2)cos^2(theta)

I = Io/4

Io/4 = Io/2 cos^2(theta)

theta = 45 degrees

(b) theta = 30 degrees

I = (Io/2) cos^2(30)

I = 0.375 Io

I/Io =0.375 =3/8